# When Less is More: Evolutionary Dynamics of Deception in a Sender-Receiver Game

### v1

@@ 1 Introduction
While behavioral economics experiments offer valuable insights into honesty and deception [11] (see [12, 13] for related evolutionary game theory analyses), they typically involve transparent payoff structures where all participants are aware of the consequences of both truthful and deceptive actions. This full information setting, while useful for studying individual behavior, may not fully capture the dynamics of disinformation, where the intent to mislead is often obscured. To better model this aspect, we consider a simplified sender-receiver game where the sender’s goal is to manipulate the receiver’s belief, and the receiver aims to infer the true state of the world. Specifically, the sender, instead of rolling a die as in the original behavioral economics experiments [11], flips a coin and reports the outcome to the receiver, who must then decide whether to believe the report. This simplification, by removing the element of chance, allows us to focus on the strategic interaction between sender and receiver and analyze the evolutionary pressures favoring truth or deception. Senders can either report the true coin flip or lie, while receivers can choose to believe or disbelieve. Importantly, in our evolutionary framework, players only have access to the payoffs of others playing the same role (sender or receiver), mimicking the information asymmetry often present in real-world disinformation scenarios.
@@
To investigate the stochastic dynamics of finite populations, we employ extensive Monte Carlo simulations combined with finite-size scaling analysis [19]. For smaller populations, these simulations are validated by numerical solutions of the corresponding birth-death process [20]. Our key finding is that the mean fixation time scales linearly with population size. This implies that the stochastic dynamics quickly transitions away from the oscillatory regime predicted by the replicator equation and rapidly converges to one of the absorbing states. This contrasts sharply with the typical exponential scaling of the mean time to escape coexistence equilibria [21, 22]. Furthermore, we identify a threshold in the payoff parameter space that separates regions where senders or receivers are more likely to achieve fixation. We find that the role with less at stake - or, equivalently, the role that evolves more slowly - is more likely to win the game. While the initially rarer strategy within the slower evolving role has a higher probability of fixation, this effect diminishes as population size increases. Surprisingly, imitation dynamics favor the fixation of the strategy with lower growth rate and initial frequency, contradicting the typical outcome of replicator dynamics for biological macromolecules [23, 24].
@@ 7 Discussion
In conclusion, our finite-population analysis of this sender-receiver game, which models the strategic interplay between deceptive senders and discerning receivers, reveals a counterintuitive outcome: the role characterized by the larger difference in payoffs between successful and unsuccessful interactions (i.e., $b_{s}+c_{s}$ for senders and $b_{r}+c_{r}$ for receivers) is ultimately more susceptible to defeat. Therefore, the role with the greater potential reward is, paradoxically, the more vulnerable.

### v2

@@ 1 Introduction
While behavioral economics experiments offer valuable insights into honesty and deception [11], they typically involve transparent payoff structures where all participants are aware of the consequences of both truthful and deceptive actions. This full information setting, while useful for studying individual behavior, may not fully capture the dynamics of disinformation, where the intent to mislead is often obscured. To better model this aspect, we consider a simplified sender-receiver game where the sender’s goal is to manipulate the receiver’s belief, and the receiver aims to infer the true state of the world. Specifically, the sender, instead of rolling a die as in the original behavioral economics experiments [11], flips a coin and reports the outcome to the receiver, who must then decide whether to believe the report. This simplification, by removing the element of chance, allows us to focus on the strategic interaction between sender and receiver and analyze the evolutionary pressures favoring truth or deception. Senders can either report the true coin flip or lie, while receivers can choose to believe or disbelieve. Importantly, in our evolutionary framework, players only have access to the payoffs of others playing the same role (sender or receiver), mimicking the information asymmetry often present in real-world disinformation scenarios.
The simplified setup of the sender-receiver game, in which the original random dice roll is replaced by a coin toss, captures the core antagonistic interaction between senders and receivers and focuses on the fundamental evolutionary dynamics of deception. This inherent antagonism is somewhat obscured by the significant random component introduced by guessing the outcome of a dice roll. In particular, a disbeliever may not receive the true outcome even when the sender is lying (the probability of guessing correctly is 1/5), while a believer always receives the correct outcome when the sender is truthful. The different types of lies (e.g., ‘black’ or ‘white’) considered in the behavioral economics experiments [11] are effectively determined by the payoff matrix, and can therefore be studied within the simplified coin toss scenario. However, in the context of deception and misinformation, ‘black’ lies – those that harm a believing receiver – are of primary concern.
Evolutionary game theory is particularly well-suited for this study because it allows us to model the interaction between senders and receivers as a dynamic, adaptive process. It accounts for the fact that senders and receivers continuously adjust their strategies in response to each other, based on their relative success. This dynamic interplay is crucial for understanding the spread and persistence of misinformation and deception. Specifically, it enables us to examine the long-term consequences of different strategies and identify stable or unstable outcomes, providing insights into the persistence or collapse of deceptive behaviors. Moreover, evolutionary game theory has a long-standing history in the study of the evolution of deception. As noted by Sober, this framework provides a powerful tool for analyzing the adaptive nature of lying [3]. Furthermore, the original behavioral experiments on the evolution of lying [11] were designed to facilitate imitation between subjects, a scenario subsequently analyzed using an evolutionary game theory formulation [12, 13]. These studies highlight the relevance and applicability of evolutionary game theory to understanding the dynamics of deception.
@@
To investigate the stochastic dynamics of finite populations, we employ extensive Monte Carlo simulations [19, 20] combined with finite-size scaling analysis [21]. Use of Monte Carlo and finite-size scaling in conjunction is a common practice in statistical physics, especially when the interest is to infer the behavior of infinitely large systems by simulating small systems [21]. For smaller populations, these simulations are validated by numerical solutions of the corresponding birth-death process [22]. Our key finding is that the mean fixation time scales linearly with population size. This implies that the stochastic dynamics quickly transitions away from the oscillatory regime predicted by the replicator equation and rapidly converges to one of the absorbing states. This contrasts sharply with the typical exponential scaling of the mean time to escape coexistence equilibria [23, 24]. Furthermore, we identify a threshold in the payoff parameter space that separates regions where senders or receivers are more likely to achieve fixation. We find that the role with less at stake - or, equivalently, the role that evolves more slowly - is more likely to win the game. While the initially rarer strategy within the slower evolving role has a higher probability of fixation, this effect diminishes as population size increases. Surprisingly, imitation dynamics favor the fixation of the strategy with lower growth rate and initial frequency, contradicting the typical outcome of replicator dynamics for biological macromolecules [25, 26].
@@ 7 Discussion
A significant advantage of both the deterministic and stochastic versions of our model is the robustness of our results to the specific values of the payoff parameters $b_{s}$, $c_{s}$, $b_{r}$ and $c_{r}$. Instead, the dynamics are governed solely by the ratio $\gamma$, defined in eq. (12). While directly mapping these simplified model parameters to precise real-world scenarios presents challenges, we can offer a clear and insightful interpretation for the ratio $\gamma$. Specifically, $\gamma>1$ signifies that receivers exhibit a higher evolutionary rate and greater stakes than senders, whereas $\gamma<1$ indicates the converse.
@@
While our analysis provides insights into the evolutionary dynamics of deception in a well-mixed population, real-world interactions are often structured by social networks [34]. Investigating the impact of network topology on our sender-receiver game represents a natural and important extension. For instance, in scale-free networks, highly connected individuals could play a disproportionate role in the spread of deceptive strategies, whereas in small-world networks, local clustering could lead to the formation of strategy clusters, influencing the overall evolutionary outcome [35]. Future research could explore these effects by incorporating network structure into our model. This, however, poses a particularly complex challenge, as it necessitates the introduction of three, in principle, distinct networks forming a multilayer network [36]. The first defines the undirected links between senders, dictating the imitation dynamics among them. The second similarly defines the undirected links between receivers, governing their imitation dynamics. Finally, the third is a directed network connecting senders to receivers, representing the flow of information.
Our study of deception in a sender-receiver game intersects with broader research on information diffusion and opinion dynamics. Indeed, evolutionary game theory has been fruitfully applied to these areas [37, 38, 39]. In particular, studies of information diffusion in structured populations highlight how factors such as public knowledge, interaction feedback, and network structure influence the spread of competing opinions or information. However, our approach differs significantly in at least two key aspects. First, agents in our model have distinct and fixed roles (senders and receivers), meaning that there is no information flow between agents with the same role, as seen in typical information diffusion models. In our model, information flows only from sender to receiver. Second, the unique asymmetric payoff structure of our model leads to cyclical behavior in the infinite population limit, in contrast to the equilibrium scenario typically observed in information diffusion models.
In conclusion, our finite-population analysis of this sender-receiver game, which models the strategic interplay between deceptive senders and discerning receivers, reveals a counterintuitive outcome: the role with the larger difference in payoffs between successful and unsuccessful interactions (i.e., $b_{s}+c_{s}$ for senders and $b_{r}+c_{r}$ for receivers) is ultimately more susceptible to defeat. This suggests that over-reliance on high-stakes deception can be counterproductive in the long run, favoring instead subtle, low-risk manipulation strategies for sustained influence. Indeed, numerous manipulative social media tactics exemplify these principles. For example, content designed for algorithmic amplification, such as highly emotional or divisive material, minimizes risk by leveraging algorithms for dissemination. Similarly, the creation of fictitious online accounts for spreading disinformation offers reduced visibility and diminishes the risk of direct exposure. Moreover, the gradual erosion of democratic norms and the normalization of extremist rhetoric are facilitated by incrementally introducing extreme or controversial concepts, beginning with less radical iterations, thereby ensuring that each step elicits a less pronounced reaction.